OpenStudy (anonymous):

a rock is thrown horizontally from the top of a water tower lands 20.0 m from the base of the tower. if the rock was initially thrown at a velocity of 10.0 m/s. how high is the water tower? what is the final velocity of the rock?

OpenStudy (anonymous):

Time = distance traveled / velocity = 20/10 = 2 s. Vertical distance traveled during 2s = 1/2gt^2 = 4.9*4 = 19.6 m Using energy principle Initial total energy = mgh + 0.5mu^2 and final total energy 0.5mv^2 Equating the two, v^2 = u^2 + 2gh = 10^2 +2* 9.8*19.6 v = 22.0 m/s